# LoRA-Contextualizing Adaptation of Large Multimodal Models for Long Document Understanding

### v1

@@ Abstract
Large multimodal models (LMMs) have recently shown great progress in text-rich image understanding, yet they still struggle with complex, multi-page visually-rich documents. Traditional methods using document parsers for retrieval-augmented generation suffer from performance and efficiency limitations, while directly presenting all pages to LMMs leads to inefficiencies, especially with lengthy ones. In this work, we present a novel framework named LoRA-Contextualizing Adaptation of Large multimodal models (LoCAL [^1] The source code is available at: https://github.com/puar-playground/LoCAL), which can broaden horizons of any LMM to support long-document understanding. We demonstrate that LMMs themselves can be an effective multimodal retriever to fetch relevant pages and then answer user questions based on these pages. LoCAL is implemented with two specific LMM adapters, one for evidence page retrieval and the other for question answering. Empirical results show state-of-the-art performance on public benchmarks, demonstrating the effectiveness of LoCAL.
@@ 1 Introduction
Documents serve as a critical medium for the preservation and dissemination of information, with millions produced annually. These documents are not limited to simple text; they encompass complex layouts and a variety of modalities such as text, tables, charts, and images. Visually-rich document understanding (VDU) is thus an essential and challenging area of research. Recently, Large Multimodal Models (LMMs) has emerged, showcasing remarkable abilities to process and understand documents. These models span both proprietary and open-source domains, like GPT-4o (OpenAI, 2023), Gemini-1.5 (Team et al., 2023), and Claude-3 among closed-source models, and InternLM-XC2-4KHD (Dong et al., 2024), InternVL-Chat (Chen et al., 2023b), LLaVA-NeXT (Liu et al., 2024a), Mini-CPM (Hu et al., 2024), mPLUG-DocOwl (Ye et al., 2023b), and TextMonkey (Liu et al., 2024d) in open-source space. Their performance has been particularly notable in single-page DU tasks demonstrated on datasets like DocVQA (Mathew et al., 2021), ChartQA (Masry et al., 2022) and InfoVQA (Mathew et al., 2022).
@@
In real-world applications, they often present documents that are much longer, containing dozens or hundreds of pages(Ma et al., 2024c; Tanaka et al., 2023; Islam et al., 2023; Zhu et al., 2021). Addressing the understanding of such lengthy documents presents LMMs with new challenges (Ma et al., 2024c). One way is to utilize a classical document parser (Rausch et al., 2021) to extract information and formulate a prompt for LLM (Wang et al., 2023; Lamott et al., 2024), which is difficult to recover the layout in prompts and suffers performance degeneration from the document parser. The other way is to exploit the long context windows of large models, allowing them to take multiple pages at once. However, most of the input pages are not relevant to user requests, and efficiency will be compromised when the document contains hundreds of pages Ma et al. (2024c); Islam et al. (2023) or there is a document collection (Tito et al., 2021).
@@
In this work, we first retrieve evidence pages to obtain relevant information within a vast and varied landscape of content. Unlike using a classical document parser, we propose using LMMs as the information encoder, which have shown great generalization ability as they have been trained on a huge text corpus. After obtaining the embedding of each page, we further utilize contextualized late interaction for relevance scoring (Khattab & Zaharia, 2020). This design shows significantly better efficiency and accuracy than using the classical document parser to extract information. Top-$k$ pages are then selected from hundreds of pages and provided to LMMs to answer user questions on documents.
@@
Figure 1: Overview of the LoCAL pipeline. The multi-page document and query are encoded by a customized LMM (yellow). The most relevant page is retrieved through similarity-based matching, and a fine-tuned LMM (blue) generates the final answer from the evidence.
@@
Based on this design demonstrated in Figure 1, we introduce the LoCAL framework for multi-page document understanding, which includes modules for evidence page retrieval and answer generation. Our contributions can be summarized as follows.
@@
  • We propose a novel framework named LoCAL to broaden the horizons of LMMs, where we use intermediate LMMs hidden embedding for efficient question-based evidence page retrieval.
@@
  • We finetune LMMs through dual LoRA adapters for evidence page retrieval and question answering, respectively, enabling LoCAL to be edge-friendly with great memory efficiency.
@@
  • We collect a visually-rich document QA dataset, LoCAL-bench, comprising nine domains including magazine, flyer, newsletter, product manual, and presentations, etc. This dataset is built upon web-crawl documents, containing 226 documents and 471 question answer pairs.
@@
  • We empirically show that LoCAL, with only 4B parameters, achieves state-of-the-art performance on LoCAL-bench and four public benchmarks, rivaling Gemini-1.5-pro on MMLongBench-Doc and demonstrating its effectiveness.
@@ 2 Related Work
Visual Document Understanding (VDU) is the field focused on interpreting text-rich images such as tables (Zhong et al., 2019), charts (Masry et al., 2022), and webpage screenshots (Liu et al., 2024c; Tanaka et al., 2021). These images are complex, featuring a mix of text and visual elements that convey abundant information (Gu et al., 2024). To evaluate multimodal document understanding, tasks range from low-level recognition tasks, such as information extraction, to high-level cognitive tasks, such as visual question answering (Mathew et al., 2020). Models in VDU are typically divided into two categories: OCR-dependent (Xu et al., 2020) and OCR-free (Kim et al., 2022), based on their reliance on Optical Character Recognition (OCR). OCR-dependent models are often trained to synchronize textual and visual data. For instance, UDoP (Tang et al., 2023) is pre-trained to restore obscured textual and layout details using both image and partial text inputs. OCR-free approaches must include text recognition training. Dount (Kim et al., 2022) is an example of an OCR-free model that focuses on producing unbroken text sequences, disregarding structural details. In contrast, Pix2Struct (Lee et al., 2023a), another OCR-free model, focuses on interpreting the structure by creating HTML DOM trees from webpage screenshots. However, this technique does not easily transfer to other image types. Our method focuses on the visual question-answering task, specifically targeting questions over long documents consisting of multiple pages of multimodal information.
@@
Augmenting language models with information from various knowledge sources has been found to boost their performance in different NLP tasks. The Dense Passage Retriever (DPR) (Karpukhin et al., 2020) trains its retrieval mechanism with documents from within the same batch, using contrastive learning with negatively sampled examples, which enhances its capabilities in open-domain question answering. Document Screenshot Embedding (DSE) (Ma et al., 2024b) uses large multimodal models (LMMs) as encoders for both document screenshots and queries, training through contrastive learning to achieve enhanced multimodal retrieval. Both REALM (Guu et al., 2020) and Retrieval-Augmented Generation (RAG) (Gao et al., 2023b) consider the passages they retrieve as hidden variables and train the retrieval and generation components together, improving the efficiency of the retrieval-augmented generation approach. Taking cues from textual RAG, the Plug-and-play (Chen et al., 2024c) approach uses GradCAM (Selvaraju et al., 2020) to fetch pertinent image segments corresponding to a given query. The MuRAG (Chen et al., 2022) model introduces a multimodal retrieval-augmented Transformer that utilizes an external multimodal memory for language generation enhancement. Unlike other approaches that retrieve information from various knowledge sources, LoCAL focuses on retrieving relevant evidence pages from a given document. This helps LMMs generate accurate and explainable answers based on the retrieved content.
Large Multimodal Models
@@
While Large Language Models (LLMs) excel at text-only question answering (QA) (Dasigi et al., 2021; Lee et al., 2023b), they cannot process other modalities. To enable multimodal tasks like Visual Question Answering (VQA), Large Multimodal Models (LMMs) transform images and videos into visual tokens that LLMs can understand. To train these LMMs, MiniGPT-4 (Zhu et al., 2023) leverages ChatGPT to produce data compliant with high-quality instructions, while LLaVA (Liu et al., 2023b) prompts GPT-4 with image captions and bounding boxes. Chen et al. (2023a; 2024a) have prompted OpenAI GPT-4V to generate more than 1M pieces of quality data to train LMMs. LLaMA-Adapter (Zhang et al., 2023; Gao et al., 2023a) and mPLUG-Owl (Ye et al., 2023b) align text and image features with large-scale image-text pairs. InstructBLIP (Dai et al., 2023) has restructured 13 vision-language tasks to fit an instruction-based approach. mPLUG-Owl (Ye et al., 2023a; b) implements multi-task instruction fine-tuning with public document datasets. Recent research (Liu et al., 2023a; 2024a; Bai et al., 2023; Dong et al., 2024; Xu et al., 2024; Luo et al., 2024) improves visual encoders by increasing resolution, leading to significant advances in downstream applications but also raising memory costs, especially in multi-page tasks. Our method addresses this by extending LMMs to handle multi-page documents, retrieving only relevant pages to reduce computation and avoid distractions from long visual token sequences.
3 LoCAL Method
@@
Multi-page document understanding aims to answer questions related to long and complex documents containing both text and images from users. We denote a document of $n$-pages as a sequence of images, $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}\}$. Text token sequence of a question $q$ is denoted as $\{q_{1},q_{2},\dots,q_{n}\}$. Traditional approaches that begin with a parsing step to extract content elements such as images, tables, and forms from documents, then generate answers based on these contents using LLMs (Saad-Falcon et al., 2023; Wang et al., 2023). Here, we first consider using LMMs to handle this task and avoiding the heuristic document parsing process, where we directly convert each page into a single image. It is not desired, as most pages in a document are irrelevant to user questions and performing an evidence page retrieval can further enhance the efficiency.
@@
We introduce LoCAL, a method that efficiently leverages the capabilities of pre-trained large multimodal models (LMMs) for long document question-answering (QA). LoCAL can broaden the horizon of LMMs to answer questions over long documents or document collections with hundreds of pages. This finding is based on the fact that hidden states of LMMs can be effective page representations for question-based retrieval, as shown in Section 5.6. This representation ability can be further enhanced with contrastive training using a LoRA adapter, demonstrating surprising retrieval performance of LMMs. Furthermore, we can finetune a LoRA-adpter of QA to further enhance the performance of LoCAL on specific domains. In summary, we first retrieve evidence pages to rank these images based on their relevance score to a given question $q$, then select the most relevant images, which are then fed into the LMM to generate the answer. In this section, we introduce the LoCAL architecture in Section 3.1, retrieval training in Section 3.2 and dual-adapter designs in Section 3.3.
@@
Figure 2: Model overview of LoCAL. It contains two modules, which are finetuned using LoRA (Hu et al., 2021), sharing the same pretrained LLM backbone. The retrieval module selects evidence pages for the other QA module, which provides responses to user questions.
@@ 3.1 Architecture
Figure 2 presents an overview of our model architecture, which comprises two LMM-based modules for the retrieval of evidence pages and question answers.
@@ 3.2 Contextualized Late Interaction
1:Pre-trained LMM $\{f_{v}$, $f_{l}^{r}\}$, training batch of evidence image and question pairs $\{(\mathbf{X}_{1},\mathbf{y}_{1}),\cdots,(\mathbf{X}_{b},\mathbf{y}_{b})\}$.
@@ 3.3 Parameter Sharing via Dual-Adapter
To reduce memory usage, we optimize the model by sharing a single LMM that includes both the vision encoder $f_{v}$ and the language model $f_{l}$ across both the retrieval and QA modules. To accommodate the different tasks required by each module, we insert two sets of adapters into the $f_{l}$ using the LoRA method (Hu et al., 2021). In the retrieval module, we use a set of adapters $\Delta\mathbf{W}_{r}$ to create the retrieval-LLM, $f_{l}^{r}$. For the QA module, a different set of adapters $\Delta\mathbf{W}_{a}$ is added to the $f_{l}$, creating the QA-LLM, $f_{l}^{r}$. In this way, we support both tasks using a single LLM and vision encoder, adding only $\sim$2additional parameters.
4 LoCAL-Bench
@@
About 4,000 PDF documents are crawled from the Web and contents of these documents are extracted via the a document parser [^2] Adobe Extract API: https://developer.adobe.com/document-services/apis/pdf-extract/. We keep the document with figures and throw away text-only or scan documents. To select documents with specific types of figures, we build a figure scheme that includes 19 figure types after reviewing different documents. We find some types of figures are not informative, such as logo and banner. We use the pretrained CLIP model ViT-L/14-336 (Radford et al., 2021) to perform a figure classification on the extracted figures and keep 6 out of 19 types of figures, including tables, maps, diagrams, infographics, data charts, workflows, and screenshots. After that, we also annotate the document types for all selected documents.
@@
Document parser returns all document elements in JSON format and the figures are saved separately as image files. We retrieve the JSON file for the document to obtain the contexts of each figure. Then we combine the figures with their contexts and use GPT-4o (API version 2024-02-15-preview) to generate QA pairs. For the GPT-4o prompts, we provide two demonstrations and ask GPT-4o to generate a QA pair. In addition, we perform automatic verification using GPT-4o to ensure the quality of the generation. Specifically, we only provide the figure to GPT-4o and ask it with the generated question; if GPT-4o can answer it correctly, we will keep that QA pair in the LoCAL Bench. This heuristic filter ensures that the answers are from document figures and double-checks the correctness of generated answers.
@@
LoCAL-Bench contains 226 documents and 471 human-verified question-answer pairs. Figure 3 shows the distributions of the document types and the length distribution by document type. LoCAL-Bench has a great diversity of documents compared to previous work (Tanaka et al., 2023; Islam et al., 2023; Ma et al., 2024c).

### v2

# SV-RAG: LoRA-Contextualizing Adaptation of MLLMs for Long Document Understanding

@@ Abstract
Multimodal large language models (MLLMs) have recently shown great progress in text-rich image understanding, yet they still struggle with complex, multi-page visually-rich documents. Traditional methods using document parsers for retrieval-augmented generation suffer from performance and efficiency limitations, while directly presenting all pages to MLLMs leads to inefficiencies, especially with lengthy ones. In this work, we present a novel framework named Self-Visual Retrieval-Augmented Generation (SV-RAG ), which can broaden horizons of any MLLM to support long-document understanding. We demonstrate that MLLMs themselves can be an effective multimodal retriever to fetch relevant pages and then answer user questions based on these pages. SV-RAG is implemented with two specific MLLM adapters, one for evidence page retrieval and the other for question answering. Empirical results show state-of-the-art performance on public benchmarks, demonstrating the effectiveness of SV-RAG.
@@ 1 Introduction
Documents serve as a critical medium for the preservation and dissemination of information, with millions produced annually. These documents are not limited to simple text; they encompass complex layouts and a variety of modalities such as text, tables, figures, and charts. Visually-rich document understanding (VDU) is thus an essential and challenging area of research. Recently, Multimodal Large Language Models (MLLMs) has emerged, showcasing remarkable abilities to process and understand documents. These models span both proprietary and open-source domains, like GPT-4o (OpenAI, 2023), Gemini-1.5 (Team et al., 2023), and Claude-3 among proprietary models, and InternLM-XC2-4KHD (Dong et al., 2024), InternVL-Chat (Chen et al., 2023b), LLaVA-NeXT (Liu et al., 2024a), Mini-CPM (Hu et al., 2024), mPLUG-DocOwl (Ye et al., 2023b), LLaVAR (Zhang et al., 2023b) and TextMonkey (Liu et al., 2024d) in open-source space. Their performance has been particularly notable in single-page DU tasks demonstrated on datasets like DocVQA (Mathew et al., 2021), ChartQA (Masry et al., 2022) and InfoVQA (Mathew et al., 2022).
@@
In real-world applications, they often present documents that are much longer, containing dozens or hundreds of pages(Ma et al., 2024d; Tanaka et al., 2023; Islam et al., 2023; Zhu et al., 2021). Addressing the understanding of such lengthy documents presents MLLMs with new challenges (Ma et al., 2024d). One way is to utilize a classical document parser (Rausch et al., 2021) to extract information and formulate a prompt for LLM (Wang et al., 2023; Lamott et al., 2024), which is difficult to recover the layout in prompts and suffers performance degeneration from the document parser. The other way is to exploit the long context windows of large models, allowing them to take multiple pages at once. However, most of the input pages are not relevant to user requests, and efficiency will be compromised when the document contains hundreds of pages Ma et al. (2024d); Islam et al. (2023) or there is a document collection (Tito et al., 2021).
@@
In this work, we first retrieve evidence pages to obtain relevant information within a vast and varied landscape of content. Unlike using a classical document parser, we propose using MLLMs as the visual retriever, which have shown great generalization ability as they have been trained on a huge text corpus. After obtaining the embedding of each page, we further utilize contextualized late interaction for relevance scoring (Khattab & Zaharia, 2020). This design shows significantly better efficiency and accuracy than using the classical document parser to extract information. Top-$k$ pages are then selected from hundreds of pages and provided to MLLMs to answer user questions on documents.
@@
Figure 1: Overview of the SV-RAG pipeline. The multi-page document and query are encoded by a customized MLLM (yellow). The most relevant page is retrieved through similarity-based matching, and a fine-tuned MLLM (blue) generates the final answer from the evidence.
@@
Based on this design demonstrated in Figure 1, we introduce the SV-RAG framework for multi-page document understanding, which includes modules for evidence page retrieval and answer generation. Our contributions can be summarized as follows.
@@
  • We propose a novel framework named SV-RAG to broaden the horizons of MLLMs, where we use intermediate MLLMs hidden embedding for efficient question-based evidence page retrieval.
@@
  • We finetune MLLMs through dual LoRA adapters for evidence page retrieval and question answering, respectively, enabling SV-RAG to be edge-friendly with great memory efficiency.
@@
  • We collect a visually-rich document QA dataset, VisR-Bench, comprising nine domains including magazine, flyer, newsletter, product manual, and presentations, etc. This dataset is built upon web-crawl documents, containing 226 documents and 471 question answer pairs.
@@
  • We empirically show that SV-RAG, with only 4B parameters, achieves state-of-the-art performance on VisR-Bench and four public benchmarks, rivaling Gemini-1.5-pro on MMLongBench-Doc and demonstrating its effectiveness.
@@ 2 Related Work
Visual Document Understanding (VDU) is the field focused on interpreting text-rich images such as tables (Zhong et al., 2019), charts (Masry et al., 2022), and webpage screenshots (Liu et al., 2024c; Tanaka et al., 2021). These images are complex, featuring a mix of text and visual elements that convey abundant information (Gu et al., 2024). To evaluate multimodal document understanding, tasks range from low-level recognition tasks, such as information extraction, to high-level cognitive tasks, such as visual question answering (Mathew et al., 2020). Models in VDU are typically divided into two categories: OCR-dependent (Xu et al., 2020) and OCR-free, based on their reliance on Optical Character Recognition (OCR). OCR-dependent models are often trained to synchronize textual and visual data. For instance, UDoP (Tang et al., 2023) is pre-trained to restore obscured textual and layout details using both image and partial text inputs. OCR-free approaches must include text recognition training. Dount (Kim et al., 2022) is an example of an OCR-free model that focuses on producing unbroken text sequences, disregarding structural details. In contrast, Pix2Struct (Lee et al., 2023a), another OCR-free model, focuses on interpreting the structure by creating HTML DOM trees from webpage screenshots. However, this technique does not easily transfer to other image types. Our method focuses on the visual question-answering task, specifically targeting questions over long documents consisting of multiple pages of multimodal information.
@@
Augmenting language models with information from various knowledge sources has been found to boost their performance in different NLP tasks. The Dense Passage Retriever (DPR) (Karpukhin et al., 2020) trains its retrieval mechanism with documents from within the same batch, using contrastive learning with negatively sampled examples, which enhances its capabilities in open-domain question answering. Document Screenshot Embedding (DSE) (Ma et al., 2024c) uses MLLMs as encoders for both document screenshots and queries, training through contrastive learning to achieve enhanced multimodal retrieval. Both REALM (Guu et al., 2020) and Retrieval-Augmented Generation (RAG) (Gao et al., 2023b) consider the passages they retrieve as hidden variables and train the retrieval and generation components together, improving the efficiency of the retrieval-augmented generation approach. Taking cues from textual RAG, the Plug-and-play (Chen et al., 2024d) approach uses GradCAM (Selvaraju et al., 2020) to fetch pertinent image segments corresponding to a given query. The MuRAG (Chen et al., 2022) model introduces a multimodal retrieval-augmented Transformer that utilizes an external multimodal memory for language generation enhancement. Unlike other approaches that retrieve information from various knowledge sources, SV-RAG focuses on retrieving relevant evidence pages from a given document. This helps MLLMs generate accurate and explainable answers based on the retrieved content. MM-GEM Ma et al. (2024a) trains an MLLM with a hybrid loss for similarity computation and caption generation on natural images. In contrast, our approach targets visually rich documents, using two LoRA modules to specialize in each task.
Multimodal Large Language Models
@@
While Large Language Models (LLMs) excel at text-only question answering (QA) (Dasigi et al., 2021; Lee et al., 2023b), they cannot process other modalities. To enable multimodal tasks like Visual Question Answering (VQA), MLLMs transform images and videos into visual tokens that LLMs can understand. To train these MLLMs, MiniGPT-4 (Zhu et al., 2023) leverages ChatGPT to produce data compliant with high-quality instructions, while LLaVA (Liu et al., 2023b) prompts GPT-4 with image captions and bounding boxes. Chen et al. (2023a; 2024a) have prompted OpenAI GPT-4V to generate more than 1M pieces of quality data to train MLLMs. LLaMA-Adapter (Zhang et al., 2023a; Gao et al., 2023a) and mPLUG-Owl (Ye et al., 2023b) align text and image features with large-scale image-text pairs. InstructBLIP (Dai et al., 2023) has restructured 13 vision-language tasks to fit an instruction-based approach. mPLUG-Owl (Ye et al., 2023a; b) implements multi-task instruction fine-tuning with public document datasets. Recent research (Liu et al., 2023a; 2024a; Bai et al., 2023; Dong et al., 2024; Xu et al., 2024; Luo et al., 2024) improves visual encoders by increasing resolution, leading to significant advances in downstream applications but also raising memory costs, especially in multi-page tasks. TextMonkey Liu et al. (2024d) compresses visual tokens using a token resampler. Our method extends MLLMs to handle multi-page documents by retrieving relevant pages, reducing computation and distractions from long token sequences.
3 SV-RAG Method
@@
Multi-page document understanding aims to answer questions related to long and complex documents containing both text and images from users. We denote a document of $n$-pages as a sequence of images, $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}\}$. Text token sequence of a question $q$ is denoted as $\{q_{1},q_{2},\dots,q_{n}\}$. Traditional approaches that begin with a parsing step to extract content elements such as images, tables, and forms from documents, then generate answers based on these contents using LLMs (Saad-Falcon et al., 2023; Wang et al., 2023). Here, we first consider using MLLMs to handle this task and avoiding the heuristic document parsing process, where we directly convert each page into a single image. It is not desired, as most pages in a document are irrelevant to user questions and performing an evidence page retrieval can further enhance the efficiency.
@@
We introduce SV-RAG, a method that efficiently leverages the capabilities of pre-trained MLLMs for long document question-answering (QA). SV-RAG can broaden the horizon of MLLMs to answer questions over long documents or document collections with hundreds of pages. This finding is based on the fact that hidden states of MLLMs can be effective page representations for question-based retrieval, as shown in Section 5.6. This representation ability can be further enhanced with contrastive training using a LoRA adapter, demonstrating surprising retrieval performance of MLLMs. Furthermore, we can finetune a LoRA-adpter of QA to further enhance the performance of SV-RAG on specific domains. In summary, we first retrieve evidence pages to rank these images based on their relevance score to a given question $q$, then select the most relevant images, which are then fed into the MLLM to generate the answer. In this section, we introduce the SV-RAG architecture in Section 3.1, retrieval training in Section 3.2 and dual-adapter designs in Section 3.3.
@@
Figure 2: Model overview of SV-RAG. It contains two modules, which are finetuned using LoRA (Hu et al., 2021), sharing the same pretrained multimodal LLM backbone. The retrieval module selects evidence pages for the other QA module, which provides responses to user questions.
@@ 3.1 Architecture
Figure 2 presents an overview of our model architecture, which comprises two MLLM-based modules for the retrieval of evidence pages and question answers.
@@ 3.2 Contextualized Late Interaction
1:Pre-trained MLLM $\{f_{v}$, $f_{l}^{r}\}$, training batch of evidence image and question pairs $\{(\mathbf{X}_{1},\mathbf{y}_{1}),\cdots,(\mathbf{X}_{b},\mathbf{y}_{b})\}$.
@@ 3.3 Parameter Sharing via Dual-Adapter
To reduce memory usage, we optimize the model by sharing a single MLLM that includes both the vision encoder $f_{v}$ and the language model $f_{l}$ across both the retrieval and QA modules. To accommodate the different tasks required by each module, we insert two sets of adapters into the $f_{l}$ using the LoRA method (Hu et al., 2021). In the retrieval module, we use a set of adapters $\Delta\mathbf{W}_{r}$ to create the retrieval-LLM, $f_{l}^{r}$. For the QA module, a different set of adapters $\Delta\mathbf{W}_{a}$ is added to the $f_{l}$, creating the QA-LLM, $f_{l}^{r}$. In this way, we support both tasks using a single LLM and vision encoder, adding only $\sim$2additional parameters.
4 VisR-Bench
@@
About 4,000 PDF documents are crawled from the Web and contents of these documents are extracted via a document parser [^1] Adobe Extract API: https://developer.adobe.com/document-services/apis/pdf-extract/. We keep the document with figures and throw away text-only or scan documents. To select documents with specific types of figures, we build a figure scheme that includes 19 figure types after reviewing different documents. We find some types of figures are not informative, such as logo and banner. We use the pretrained CLIP model ViT-L/14-336 (Radford et al., 2021) to perform a figure classification on the extracted figures and keep 6 out of 19 types of figures, including tables, maps, diagrams, infographics, data charts, workflows, and screenshots. After that, we also annotate the document types for all selected documents.
@@
Document parser returns all document elements in JSON format and the figures are saved separately as image files. We retrieve the JSON file for the document to obtain the contexts of each figure. Then we combine the figures with their contexts and use GPT-4o (API version 2024-02-15-preview) to generate QA pairs. For the GPT-4o prompts, we provide two demonstrations and ask GPT-4o to generate a QA pair. In addition, we perform automatic verification using GPT-4o to ensure the quality of the generation. Specifically, we only provide the figure to GPT-4o and ask it with the generated question; if GPT-4o can answer it correctly, we will keep that QA pair in the SV-RAG Bench. This heuristic filter ensures that the answers are from document figures and double-checks the correctness of generated answers.
@@
VisR-Bench contains 226 documents and 471 human-verified question-answer pairs. Figure 3 shows the distributions of the document types and the length distribution by document type. VisR-Bench has a great diversity of documents compared to previous work (Tanaka et al., 2023; Islam et al., 2023; Ma et al., 2024d).